2zy/y2=14/y i wan to understand step by step but its hard for me to follow!
Answer by stanbon(75887)

2zy/y2=14/y
---
Reduce the fraction on the left side:
---
[2z/y] = [14/y]
---
2z = 14
z = 7
---
Solve for "y":
[2*7/y] = [14/y]
---
y is any Real number other than zero.
=====================================
Cheers,
Stan H.